1. An electric motor in a radio-controlled car does work to accelerate the car from rest to a speed of 2.8 m/s in a timespan of 0.35 s.  If the mass of the RC car is 1.8 kg, calculate the average power delivered to the car to accelerate it.
1
Expert's answer
2021-05-09T13:07:32-0400
"F=\frac{mv}{t}\\P=\frac{mv^2}{2t}\\\\P=\frac{(1.8)(2.8)^2}{2(0.35)}=20\ W"